The Tax Formula for Individuals, A Brief Overview of Capital Gains and Losses (LO 1.3, L.O. 1.8)

In 2018, Lou has a salary of $53,300 from her job. She also has interest income of $1,600 and dividend income of $400. Lou is single and has no dependents. During the year, Lou sold silver coins held as an investment for a $7,000 loss.
Answer #1
  • Maximum permissible capital loss deduction = $ 3,000
  • Adjusted Gross Income = Salary + interest income + dividend income - maximum permissible capital loss deduction = 53,300 + 1,600 + 400 - 3,000 =   52,300   capital loss deduction is restricted to $3000
  • Standard Deduction for single in the year 2018= $  12,000   
  • Exemption = 0   
  • Taxable income = 52,300 - 12,000 = $ 39,700